OpenStudy (anonymous):

Please help asap! Medal and fan! :) 1. Use the zero product property to find the solutions to the equation x2 – 13x + 30 = 0. a) x = –10 or x = –3 b) x = –10 or x = 3 c) x = –3 or x = 10 d) x = 3 or x = 10 2. The profit a company makes from producing x tabletops is modeled by the equation P(x) = 480x – 2x2. For what number of tabletops does the company make a profit of $0? a) 100 tabletops b) 120 tabletops c) 240 tabletops d) 480 tabletops

OpenStudy (anonymous):

1. A (I factored it instead, but you end up with the same answer)

OpenStudy (anonymous):

I can't understand the equation to the second problem though. Are you trying to say:\[P(x)=480x-2x^2\]?

OpenStudy (anonymous):

Medals and fans accepted :)

OpenStudy (chris911):

X^2 + 13X + 30 = 0 Since a = 1, you need only concern yourself with b = 13 & c = 30. 30 has several factors: 1*30 = 30, 2*15 = 30, 3*10 = 30, & 6*5 = 30. But, only 3 + 10 = 13 So we have: (X + 3)(X + 10) = 0 Either X + 3 = 0 We subtract 3 from both sides of the equation. We get: X + 3 -- 3 = 0 -- 3 Simplifying yields: X = -- 3 OR X + 10 = 0 We subtract 10 from both sides of the equation. We get: X + 10 -- 10 = 0 -- 10 Simplifying yields: X = -- 10 Therefore, X = -- 3 or -- 10.

OpenStudy (anonymous):

Yes, I was meaning that. :) and thanks for the help.

OpenStudy (anonymous):

No problem! 2. C

OpenStudy (chris911):

2 would be C I think because u take 480 divided by 2

OpenStudy (anonymous):

In order to get C, solve the equation: 180x-2x^2=0

OpenStudy (anonymous):

@Chris911 , thats not the way to solve for C

OpenStudy (chris911):

great work @severus4ever ;)

OpenStudy (anonymous):

Thank you both! :)

OpenStudy (anonymous):

@Chris911 , thank you! You're explanation for Number 1 was more valid than mine.

OpenStudy (anonymous):

No problem :)

OpenStudy (chris911):

:)